The number of real solutions of the equation 3-|x| - 2|x| = 0

Grade:12th pass

1 Answers

Arun
5 years ago
Dear Manish
 
If x>= 0
1/3^|x|  - 2^|x| = 0
1 - 6^|x| = 0
x = 0
Now if X
3^|x| - 1/2^|x| = 0
6^|x| - 1 = 0
x = 0
Hence 
Only one solution exist i.e. x = 0